OpenStudy (anonymous):

Im stuck on this combine negative thing. can someone help me? 11w-9-7w=15

OpenStudy (anonymous):

6

OpenStudy (anonymous):

w=6

OpenStudy (anonymous):

could you help with the steps?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

add 9 to 15 to get 24 then take 11w and combine it with 7w and you will get 4w=24 divide 24 by 4 to get w=6

OpenStudy (jango_in_dtown):

11w-9-7w=15 implies (11w-7w)=15+9 implies 4w=24 and hence

OpenStudy (anonymous):



OpenStudy (anonymous):

Thanks for all of your help!